### v3

# Examining Multilingual Embedding Models Cross-Lingually Through LLM-Generated Adversarial Examples

Abstract
The evaluation of cross-lingual semantic search capabilities of models is often limited to existing datasets from tasks such as information retrieval and semantic textual similarity. To allow for domain-specific evaluation, we introduce Cross Lingual Semantic Discrimination (CLSD), a novel cross-lingual semantic search task that does not require a large evaluation corpus, only parallel sentences of the language pair of interest within the target domain. This task focuses on the ability of a model to cross-lingually rank the true parallel sentence higher than challenging distractors generated by a large language model. We create a case study of our introduced CLSD task for the language pair German-French in the news domain. Within this case study, we find that models that are also fine-tuned for retrieval tasks benefit from pivoting through English, while bitext mining models perform best directly cross-lingually. A fine-grained similarity analysis enabled by our distractor generation strategy indicate that different embedding models are sensitive to different types of perturbations.
Examining Multilingual Embedding Models Cross-Lingually Through LLM-Generated Adversarial Examples
Andrianos Michail, Simon Clematide, Rico Sennrich University of Zurich {andrianos.michail,simon.clematide,sennrich}@cl.uzh.ch
1 Introduction
When considering which model to use to perform semantic search within a specific domain and language pair, standard benchmarks such as MTEB (Muennighoff et al., 2023) do not cover every domain within all language pairs. We posit that one should aim to directly evaluate semantic search for the target language pair within the domain of interest but the lack of dataset for such evaluation cases makes it difficult to take fundamental decisions for the application uses, such as the choice of an embedding model or whether to use English as a pivot language to perform the task.
We propose a new simple and efficient dataset generation and evaluation scenario that simulates real-world challenges of cross-lingual semantic retrieval in large text collections. Our approach involves the use of LLMs to create adversarial examples from parallel sentences that challenge the cross-lingual embedding capabilities of multilingual models. Concretely speaking, four distractor sentences are generated for each parallel target sentence. We require these sentences to be very similar to the original target sentence in terms of syntactic structure and surface word forms, but to be semantically dissimilar. Our proposed task, Cross-Lingual Semantic Discrimination (CLSD), involves identifying the correct target sentence from among four distractor sentences in the target language, given an original sentence in the source language.
To better understand how the introduced changes in the distractors influence the semantic similarities cross- and monolingually, we perform a fine-grained analysis and correlate the semantic changes with linguistic properties at the level of part-of-speech tags.
Contributions
1. Proposing a new task that emulates cross-lingual semantic search within large text collections, requiring only a set of parallel sentences for dataset creation. 2. Publishing four such datasets within the news domain for the language pair German-French. 3. Demonstrate that within the task, certain embedding models perform better in direct cross-lingual retrieval, while others benefit from using English as a pivot language. 4. Providing insights into the cross- and monolingual semantic representation of multilingual models by a linguistically informed analysis.
2 Related Work
Original Sentence Distractor Sentence 1 Distractor Sentence 2
Die Linkspartei beschließt in Bonn ihr Programm zur Europawahl.
(The Left Party adopts its program for the European elections in Bonn.)
Die Linkspartei beschließt in Bonn ihr Programm zur Bundestagswahl.
(The Left Party adopts its program for the Bundestag elections in Bonn.)
Die Linkspartei verweigert in Bonn ihr Programm zur Europawahl.
(The Left Party refuses its program for the European elections in Bonn.)
Là-bas, la perspective de la fin de la guerre reste toujours très éloignée.
(Over there, the prospect of the end of the war is still a long way off.)
Là-bas, la fin de la guerre semble toujours très incertaine.
(Over there, the end of the war still seems very uncertain.)
Là-bas, la perspective de la paix semble toujours très éloignée.
(Over there, the prospect of peace still seems a long way off.)
Table 1: Examples of generated distractor sentences in German and French with English translations in italics. The red font marks modified text.
Cross-Lingual Semantic Search
Cross-lingual semantic search is an umbrella term for tasks that involve searching for texts with similar or relevant semantics in another language. Examples of such cross-lingual tasks are information retrieval (CLIR) (Lawrie et al., 2023, 2024), question answering (CLQA) (Lewis et al., 2020), semantic text similarity (X-STS) (Cer et al., 2017), and bitext mining (Research, ; Zweigenbaum et al., 2017). As shown in MTEB (Muennighoff et al., 2023), multilingual embedding models vary in their performance for different cross-lingual semantic search tasks and language pairs. When selecting models for a use case, we often rely on results from other domains or language pairs, as direct evaluation results are not always available.
Distractor Generation
PAWS-X (Yang et al., 2019; Zhang et al., 2019) generated adversarial examples by rule-based and backtranslation methods to measure the limitations of paraphrase identification models. More recently, XSim++ (Chen et al., 2023) used rule-based text augmentation to generate synthetic examples for tuning LASER models (replacements by antonyms, swapping entities). Within semantic search, InPars (Bonifacio et al., 2022) use solely in-domain synthetic data to finetune retrieval models that surpass strong baselines.
Model (Hugging Face Name) X-Ling. Aligned Paraph. Aligned Retrieval Aligned WMT19-DE->FR WMT21-DE->FR WMT19-FR->DE WMT21-FR->DE     Mean
Cross-Lingual CLSD Evaluation
multilingual-e5-base 91.51 86.34 88.46 81.97     87.07
multilingual-e5-large 94.43 91.38 91.45 87.57     91.21
gte-multilingual-base 90.22 90.48 89.55 91.60     90.46
paraphrase-multilingual-mpnet-base 91.31 91.15 91.11 92.95     91.63
sentence-transformers/LaBSE 95.18 94.06 94.30 94.18     94.43
Machine Translation (M2M 1.2B (Fan et al., 2021)) to EN -> CLSD Evaluation
multilingual-e5-base 90.50 89.81 89.55 93.39     90.81
multilingual-e5-large 90.22 91.83 89.95 92.95     91.24
gte-multilingual-base 89.68 90.26 89.07 92.95     90.49
paraphrase-multilingual-mpnet-base 87.92 88.13 87.64 91.15     88.71
sentence-transformers/LaBSE 90.84 91.83 90.16 94.40     91.81
Table 2: Precision@1 on the Cross-Lingual Semantic Discrimination Datasets.
3 Experiments
3.1 Cross-Lingual Semantic Discrimination (CLSD)
We propose a task that aims to evaluate multilingual embedding methods based on their capacity to find the closest cross-lingual semantic match amidst large bilingual text collections. The goal is to measure the model’s ability to identify a parallel sentence as the most similar to the original sentence in the source language. To accomplish this task, a dataset of parallel sentences is required that will be enriched with $N$ distractor sentences in the target language.
Distractors Generation Prompt Can you provide me with four tricky sentences (numbered) that look structurally and lexically similar but don’t have the same meaning. The sentences should be within similar topics and share commonalities with the original. Answer in {French/German}!
{Original Sentence in French/German}
Figure 1: GPT-4 prompt template used for the generation of distractors
In this paper, we test the hypothesis that LLMs can generate difficult examples that look similar on the surface, but convey different meanings. We prompt GPT-4 (gpt-4-0613) (OpenAI, 2023) to monolingually generate four distractors for each original sentence (see Figure 1). The distractors are requested to be structurally and lexically similar, whilst being semantically dissimilar. For each language, we manually examine 200 distractors of the original sentences and find that over 98% of the distractors meet these criteria. Example sentences are shown in Table 1.
CLSD Datasets
We generate and validate (samples of) four datasets, consisting of both language directions of German/French for the WMT19 (Barrault et al., 2019) and WMT21 (Akhbardeh et al., 2021) test sets. [^1] The data is available at link. In the 18,928 generated distractors, we observe that the LLM used different modification strategies such as single polarity inversion, entity replacement or even multiple swaps with word-level Jaccard similarity between originals and distractors at ($\mu=0.47$, $\sigma=0.17$). Table 4 illustrates samples of distractors sets.
CLSD Evaluation
To evaluate a multilingual embedding model $E$, we assess its ability to discriminate between original parallel sentence pairs and challenging distractors. For a given sentence $S$ in the source language and its corresponding translation $T$ in the target language, along with a set of $N$ distractor sentences $D$ in the target language, the model must produce an embedding of $S$ that is more similar to the embedding of $T$ than to that of any distractor $d_{i}$ in $D$.
To measure performance of embedding models, we use Precision@1 (P@1), defined as the proportion of samples in which the model ranks the true translation as more similar to the source sentence than the four distractors. For example, in Table 2, the column WMT19-DE->FR reports the P@1 of identifying the original French target sentence as more similar than its French distractors.
Models Evaluated
We examine the following five embedding models available through SentenceTransformers (Reimers and Gurevych, 2019).
multilingual-e5-base/-large (M-E5-b/l) Wang et al. (2024, 2024) transform XLM-RoBERTa (Conneau et al., 2020) into a bi-encoder through weakly supervised contrastive pre-training with multilingual data and a second step of supervised fine-tuning on retrieval datasets.
gte-multilingual-base (M-GTE) Zhang et al. (2024) pre-train a multilingual long-context encoder (8192) and derive a bi-encoder by a process similar to multilingual-e5 that also includes contrastive training against hard negatives.
paraphrase-multilingual-mpnet-base-v2 (M-MPNet) Reimers and Gurevych (2020) use knowledge distillation through cosine loss training on parallel sentences between an English paraphrase trained teacher model (paraphrase-mpnet-base-v2) and a multilingual student model (XLM-RoBERTa (Conneau et al., 2020)) to teach the student to represent sentences closely across multiple languages.
LaBSE Language-agnostic BERT Sentence Embedding (Feng et al., 2022) is trained with translation ranking loss and negative samples.
3.2 CLSD Results
The results of the models on the CLSD datasets are shown in Table 2. Regarding the direct cross-lingual evaluation, the Bitext Mining specialist LaBSE is the best model in this task with an average score of 94.43 (P@1), demonstrating its consistent ability to cross-lingually rank exact parallel text higher than the distractors. Another pattern that emerges is that models (M-E5-b/l, M-GTE) that were also trained on English-centric retrieval datasets, such as the MS Macro (Nguyen et al., 2016), perform worse within our stricter task.
In the lower half of the table, we evaluate the models using English as a pivot language on both sides, employing the M2M 1.2B translation model (Fan et al., 2021). Both LaBSE and M-MPNet show a performance decrease of 2.6/2.8 points compared to direct cross-lingual evaluation. In contrast, models with weaker performance in direct cross-lingual show only minor improvements, except the M-E5-b model, which gains 3.7 points.
We conduct a qualitative analysis of the 279 samples where LaBSE succeeds in the direct cross-lingual evaluation but fails in the pivot-through-English evaluation. Our analysis reveals that, in most cases, language-specific terms are translated into broader and common English words, reducing the distinction between the original texts and the distractors. Less frequently, we observe omitted details in the translations, hallucinations, and, rarely, poor quality translations.
3.3 Fine-grained similarity analysis
[…]
$\frac{(1-\cos(E(S),E(T)))-(1-\cos(E(S),E(d))}{\|-\nparallel_{\text{ DE }\leftrightarrow\text{ FR}}}$
Figure 2 presents model results, showing that the M-MPNet model experiences the largest overall shifts in cosine similarity when single-word modifications are made. The column “ANY” reflects changes across all items, facilitating within-model comparisons to determine which linguistically characterized subgroups exhibit more pronounced changes. Notably, LaBSE, the best performing model within CLSD, represents these single-word changes with relatively uniform similarity shifts across linguistic categories, unlike the other models. Additionally, we observe that replacing proper nouns consistently decreases semantic similarity more than replacing other parts of speech across all languages and models.
Figure 2: Cross-lingual cosine similarity change in original-distractor sentence pairs with exactly one token swapped, grouped by the PoS of the differing token.
Mono- and Cross-Lingual Change Correlations
How language-agnostic are the base models when dealing with distractors? Does it matter whether the source sentence is presented in the source or target language? We measure the monolingual similarity change of the original to the distractor by substituting $E(S)$ with $E(T)$ in the calculation. The corresponding monolingual semantic change plots can be found in Figure 3 in the Appendix. By correlating the results of the cross- and monolingual change for all one-word change cases (i.e. “ANY”): The correlation between monolingual and cross-lingual changes is strong for M-MPNet with 0.93 closely followed by M-GTE with 0.85 and LaBSE with 0.82. The M-E5-b/l shows lower correlation with 0.79 and 0.71 respectively. However, specific nuances depend on the language, model, or type of replaced word. For example, M-GTE shows a low correlation of 0.64 for proper nouns between cross- and monolingual evaluations in French, compared to 0.77 for German.
4 Conclusion
This paper contributes to cross-lingual semantic search by identifying gaps in current evaluation practices and introducing adversarial examples to test multilingual embeddings. Our research, on the German-French case study, introduces four new adversarial cross-lingual evaluation datasets for the news domain. Our comparative evaluation of direct cross-lingual and pivot through English evaluation highlights the advantage of each method based on the embedding model used. The fine-grained analyses demonstrate how generated distractors offer insights into the cross-lingual capabilities of multilingual models and detail how specific small perturbations in sentences can change these models’ produced semantic representations. Future work should extend these approaches to other language pairs and refine adversarial techniques to improve semantic search model evaluations.
Limitations
We introduce a pipeline that allows cross-lingual evaluation of semantic search between any two languages that have an available set of parallel texts. However, we evaluate only a single language pair (of interest) and only a single model for distractor generation, whereas in reality much cheaper methods might suffice. Automatic prediction of PoS tags and named entities in our analysis can result in introducing noise within the fine-grained analysis results. The performance of the stanza models and our added manual results validation minimises this risk. However, the capabilities of current models in the language of interest should be critically assessed. Furthermore, it is important to note that the examination of only the few models is due to space constraints; however, it is worth considering that valuable insights can also be gained from smaller models. We hope that larger future work could address these limitations and explore this approach more broadly, creating a large scale cross-lingual semantic search dataset parallel in many, including low-resource languages.
Acknowledgments
This research is conducted under the project Impresso – Media Monitoring of the Past II Beyond Borders: Connecting Historical Newspapers and Radio. Impresso is a research project funded by the Swiss National Science Foundation (SNSF 213585) and the Luxembourg National Research Fund (17498891).
Appendix A Appendix
Model Parameter Value
Temperature 1.0
Top-P 1.0
Table 3: Parameters used for GPT-4 via chat completion for the distractor generation. Undefined parameters are the default.
Figure 3: Monolingual cosine similarity change in original-distractor sentence pairs with exactly one token swapped, grouped by the part of speech of the differing token.
Distractor Intra-Distractor Jaccard Sim.
Original: PARIS (Reuters) - Die New Yorker Börse schloss am Freitag mit 0,68 % im Minus und beendete die Woche mit einem allgemeinen Rückgang, da sie von enttäuschenden Ergebnissen, einem Wiederanstieg der Coronavirus-Infektionen und geopolitischen Unsicherheiten beeinträchtigt wurde. /
Adv1: LONDON (Reuters) - Die Londoner Börse schloss am Dienstag mit 0,68 % im Plus und startete die Woche mit allgemeinem Wachstum, da sie von erfreulichen Ergebnissen, einem kontinuierlichen Rückgang der Coronavirus-Infektionen und geopolitischen Stabilitäten angetrieben wurde. 0.434
Adv2: BERLIN (Reuters) - Die Berliner Börse schloss am Donnerstag mit 0,68 % im Minus und beendete die Woche mit einem insgesamt flachen Verlauf, das durch gemischte Ergebnisse, eine Stagnation der Coronavirus-Infektionen und geopolitische Spannungen beeinflusst wurde. 0.300
Adv3: TOKYO (Reuters) - Die Tokioter Börse schloss am Mittwoch mit 0,68 % im Plus und begann den Monat mit allgemeiner Erholung, da sie von positiven Ergebnissen, einem eutlichen Rückgang der Coronavirus-Infektionen und geopolitischer Sicherheit profitierte. 0.359
Adv4: MADRID (Reuters) - Die Madrider Börse schloss am Montag mit 0,68 % im Minus und startete die Woche mit allgemeinem Rückgang, da sie von enttäuschenden Ergebnissen, einem Wiederanstieg der Coronavirus-Infektionen und geopolitischen Unsicherheiten beeinträchtigt wurde. 0.438
Original: Der Nasdaq verzeichnete die schlechteste Woche der letzten vier. /
Adv1: Der Nasdaq hat die beste Woche der letzten vier verzeichnet. 0.630
Adv2: Der Nasdaq verzeichnete die aktivste Woche der letzten vier. 0.623
Adv3: Der Nasdaq hat die ruhigste Woche der letzten vier verzeichnet. 0.630
Adv4: Der Nasdaq verzeichnete die turbulenteste Woche der letzten vier. 0.623
Original: Die Beamten werden im Haushalt für 2021 jedoch nicht vergessen werden. /
Adv1: Die Beamten werden im Haushalt für 2021 jedoch nicht besprochen werden. 0.818
Adv2: Die Beamten werden im Haushalt für 2021 jedoch nicht entlassen werden. 0.818
Adv3: Die Beamten werden im Haushalt für 2021 jedoch nicht befördert werden. 0.818
Adv4: Die Beamten werden im Haushalt für 2021 jedoch nicht belastet werden. 0.818
Table 4: Original-Distractor sets with their Intra-Distractor Jaccard Similarity. The red font indicates modified text.
Original Distractor Tricked Model
WMT19- DE->FR-CLSD
Kipping au congrès de die Linke sur l’Europe : l’Europe est depuis longtemps un continent d’immigration. Kipping au congrès de die Linke sur l’Asie : l’Asie est depuis longtemps un continent de diversité. MPNet-M
Par exemple, nous allons à présent recevoir un quartier général pour les munitions de l’UE, à Bruxelles, ce que les Britanniques ont jusqu’ici toujours empêché. Par exemple, nous prévoyons de lancer un bureau principal pour les ressources de l’UE, à Bruxelles, ce que les Français ont toujours refusé jusqu’ici. MPNet-M
L’appel de Macron: l’Europe est davantage qu’un "projet" L’appel de Macron: l’Europe est plus qu’une "idée". M-E5-b
Trois commissions ont accepté ma candidature, l’une d’elle était de justesse contre. Trois comités ont rejeté ma candidature, l’un d’eux était presque en faveur. M-E5-b
WMT19- FR->DE-CLSD
Würden sie dies tun, gäbe es niemanden in Europa, der sich dagegen wehren würde. Würden sie dies tun, gäbe es niemanden in Europa, der sich dafür aussprechen würde. MPNet-M M-E5-b
Um dem etwas entgegenzusetzen, ist die EU-Kommission auf Kooperation mit den Internetriesen angewiesen. Um das zu fördern, ist die EU-Kommission auf Zusammenarbeit mit den Internetgiganten angewiesen. MPNet-M M-E5-b
Und in der Sozialpolitik sehe ich bei ihr ebenfalls keinen Rechtsschwenk. Und bei ihr sehe ich in der Sozialpolitik ebenfalls keinen Linksschwenk. MPNet-M
Bohnentopf, Hähnchen mit Pommes oder Kalbfleisch mit Soße und Kartoffeln: In der Mensa-Eingangshalle der Universität Karlsruhe drängeln sich um die Mittagszeit die jungen Leute. Pizzaschnitten, Hähnchen mit Reis oder Rinderbraten mit Gemüse und Kartoffeln: Um die Abendstunde füllt sich der Festsaal der Universität Stuttgart mit hungrigen Studenten. MPNet-M
Nein. Zurückweisungen gibt es ja auch in anderen EU-Ländern. Nein, Zuwanderungen gibt es ja auch in anderen EU-Ländern. M-E5-b
Und es ist zugleich die Emanzipation der Linken von ihrem Ex-Vorsitzenden Oskar Lafontaine. Und es ist zugleich die Spaltung der Linken wegen ihrem Ex-Vorsitzenden Oskar Lafontaine. M-E5-b
Table 5: Randomly selected examples of successful distractors. The red font indicates modified text.

### v4

@@
 Abstract
-The evaluation of cross-lingual semantic search capabilities of models is often limited to existing datasets from tasks such as information retrieval and semantic textual similarity.
-To allow for domain-specific evaluation, we introduce Cross Lingual Semantic Discrimination (CLSD), a novel cross-lingual semantic search task that does not require a large evaluation corpus, only parallel sentences of the language pair of interest within the target domain.
-This task focuses on the ability of a model to cross-lingually rank the true parallel sentence higher than challenging distractors generated by a large language model.
-We create a case study of our introduced CLSD task for the language pair German-French in the news domain.
-Within this case study, we find that models that are also fine-tuned for retrieval tasks benefit from pivoting through English, while bitext mining models perform best directly cross-lingually.
-A fine-grained similarity analysis enabled by our distractor generation strategy indicate that different embedding models are sensitive to different types of perturbations.
+The evaluation of cross-lingual semantic search models is often limited to existing datasets from tasks such as information retrieval and semantic textual similarity.
+We introduce Cross-Lingual Semantic Discrimination (CLSD), a lightweight evaluation task that requires only parallel sentences and a Large Language Model (LLM) to generate adversarial distractors.
+CLSD measures an embedding model’s ability to rank the true parallel sentence above semantically misleading but lexically similar alternatives.
+As a case study, we construct CLSD datasets for German–French in the news domain.
+Our experiments show that models fine-tuned for retrieval tasks benefit from pivoting through English, whereas bitext mining models perform best in direct cross-lingual settings.
+A fine-grained similarity analysis further reveals that embedding models differ in their sensitivity to linguistic perturbations.
+[^1] We release our code and datasets under AGPL-3.0: https://github.com/impresso/cross_lingual_semantic_discrimination
 Examining Multilingual Embedding Models Cross-Lingually Through LLM-Generated Adversarial Examples
-Andrianos Michail, Simon Clematide, Rico Sennrich University of Zurich {andrianos.michail,simon.clematide,sennrich}@cl.uzh.ch
+Andrianos Michail    Simon Clematide    Rico Sennrich University of Zurich {andrianos.michail,simon.clematide,sennrich}@cl.uzh.ch
 1 Introduction
-When considering which model to use to perform semantic search within a specific domain and language pair, standard benchmarks such as MTEB (Muennighoff et al., 2023) do not cover every domain within all language pairs.
-We posit that one should aim to directly evaluate semantic search for the target language pair within the domain of interest but the lack of dataset for such evaluation cases makes it difficult to take fundamental decisions for the application uses, such as the choice of an embedding model or whether to use English as a pivot language to perform the task.
-We propose a new simple and efficient dataset generation and evaluation scenario that simulates real-world challenges of cross-lingual semantic retrieval in large text collections.
-Our approach involves the use of LLMs to create adversarial examples from parallel sentences that challenge the cross-lingual embedding capabilities of multilingual models.
+When selecting a model for semantic search in a specific domain and language pair, standard embedding benchmarks such as (Multilingual) MTEB (Muennighoff et al., 2023; Enevoldsen et al., 2025) provide only partial coverage.
+We argue that evaluation should be carried out directly on the target language pair and domain of interest.
+However, the lack of suitable datasets makes it difficult to reach well-founded decisions about applications, such as choosing an embedding model or deciding whether to use English as a pivot language.
+We propose a simple and efficient method for dataset generation and evaluation that simulates real-world challenges of cross-lingual semantic retrieval in large text collections.
+Our approach uses LLMs to create adversarial examples from parallel sentences, designed to test the cross-lingual embedding capabilities of multilingual models.
 Concretely speaking, four distractor sentences are generated for each parallel target sentence.
@@
 Our proposed task, Cross-Lingual Semantic Discrimination (CLSD), involves identifying the correct target sentence from among four distractor sentences in the target language, given an original sentence in the source language.
-To better understand how the introduced changes in the distractors influence the semantic similarities cross- and monolingually, we perform a fine-grained analysis and correlate the semantic changes with linguistic properties at the level of part-of-speech tags.
+To better understand how these distractor modifications influence semantic similarity both cross-lingually and monolingually, we perform a fine-grained analysis and correlate the semantic changes with linguistic properties at the level of part-of-speech tags.
 Contributions
-1.
-Proposing a new task that emulates cross-lingual semantic search within large text collections, requiring only a set of parallel sentences for dataset creation.
-2.
-Publishing four such datasets within the news domain for the language pair German-French.
-3.
-Demonstrate that within the task, certain embedding models perform better in direct cross-lingual retrieval, while others benefit from using English as a pivot language.
-4.
-Providing insights into the cross- and monolingual semantic representation of multilingual models by a linguistically informed analysis.
+  • We propose a new task that emulates cross-lingual semantic search within large text collections, requiring only a set of parallel sentences for dataset creation.
+  • We publish four such datasets in the news domain for the German–French language pair.
+  • We show that some embedding models perform better in direct cross-lingual retrieval, while others benefit from using English as a pivot language.
+  • We provide a linguistically informed analysis that offers insights into cross- and monolingual semantic representations of multilingual models.
 2 Related Work
-Original Sentence Distractor Sentence 1 Distractor Sentence 2
+Original Sentence $S$ Distractor Sentence 1 $D_{1}$ Distractor Sentence 2 $D_{2}$
 Die Linkspartei beschließt in Bonn ihr Programm zur Europawahl.
@@
 (Over there, the prospect of peace still seems a long way off.)
-Table 1: Examples of generated distractor sentences in German and French with English translations in italics. The red font marks modified text.
+Table 1: Examples of generated distractor sentences in German and French, with English translations provided in italics. Text in red indicates the modified segments.
 Cross-Lingual Semantic Search
@@
 Examples of such cross-lingual tasks are information retrieval (CLIR) (Lawrie et al., 2023, 2024), question answering (CLQA) (Lewis et al., 2020), semantic text similarity (X-STS) (Cer et al., 2017), and bitext mining (Research, ; Zweigenbaum et al., 2017).
-As shown in MTEB (Muennighoff et al., 2023), multilingual embedding models vary in their performance for different cross-lingual semantic search tasks and language pairs.
+As showcased in MMTEB (Enevoldsen et al., 2025), multilingual embedding models vary in their performance for different cross-lingual semantic search tasks and language pairs.
 When selecting models for a use case, we often rely on results from other domains or language pairs, as direct evaluation results are not always available.
@@
 More recently, XSim++ (Chen et al., 2023) used rule-based text augmentation to generate synthetic examples for tuning LASER models (replacements by antonyms, swapping entities).
-Within semantic search, InPars (Bonifacio et al., 2022) use solely in-domain synthetic data to finetune retrieval models that surpass strong baselines.
-Model (Hugging Face Name) X-Ling. Aligned Paraph. Aligned Retrieval Aligned WMT19-DE->FR WMT21-DE->FR WMT19-FR->DE WMT21-FR->DE     Mean
+Within semantic search, InPars (Bonifacio et al., 2022) relies solely on in-domain synthetic data to fine-tune retrieval models that surpass strong baselines.
+Lastly, Li et al.
+(2025) used AMR (Banarescu et al., 2013) to craft hard negatives and examine the robustness of English embedding models.
+Bitext Mining
+Bitext Mining is an established and most similar cross-lingual semantic search task to our proposed CLSD task.
+Bitext mining (Research, ; Zweigenbaum et al., 2017) challenges embedding models to identify an exact parallel sentence from a large pool—often consisting of thousands of unrelated or loosely related sentences-originating from large, homogeneous corpora.
+Model (Hugging Face Name) Alignment     WMT19 WMT21     Average
+Cross-
+lingual
+Para-
+phrase
+Retrie-
+val
+[…]
 Cross-Lingual CLSD Evaluation
-multilingual-e5-base 91.51 86.34 88.46 81.97     87.07
-multilingual-e5-large 94.43 91.38 91.45 87.57     91.21
-gte-multilingual-base 90.22 90.48 89.55 91.60     90.46
-paraphrase-multilingual-mpnet-base 91.31 91.15 91.11 92.95     91.63
-sentence-transformers/LaBSE 95.18 94.06 94.30 94.18     94.43
-Machine Translation (M2M 1.2B (Fan et al., 2021)) to EN -> CLSD Evaluation
-multilingual-e5-base 90.50 89.81 89.55 93.39     90.81
-multilingual-e5-large 90.22 91.83 89.95 92.95     91.24
-gte-multilingual-base 89.68 90.26 89.07 92.95     90.49
-paraphrase-multilingual-mpnet-base 87.92 88.13 87.64 91.15     88.71
-sentence-transformers/LaBSE 90.84 91.83 90.16 94.40     91.81
-Table 2: Precision@1 on the Cross-Lingual Semantic Discrimination Datasets.
+multilingual-e5-base ✓     91.51 88.46 86.34 81.97     87.07
+multilingual-e5-large ✓     94.50 91.51 91.38 87.57     91.24
+gte-multilingual-base ✓     90.22 89.68 90.48 91.60     90.50
+paraphrase-multilingual-mpnet-base ✗     91.31 91.11 91.15 92.95     91.63
+sentence-transformers/LaBSE ✗     95.18 94.30 94.06 94.18     94.43
+English Pivot Translation by M2M 1.2B (Fan et al., 2021) CLSD Evaluation
+multilingual-e5-base ✓     90.50 89.55 89.81 93.39     90.81
+multilingual-e5-large ✓     90.22 89.95 91.83 92.95     91.24
+gte-multilingual-base ✓     89.68 89.07 90.26 92.95     90.49
+paraphrase-multilingual-mpnet-base ✗     87.92 87.64 88.13 91.15     88.71
+sentence-transformers/LaBSE ✗     90.84 90.16 91.83 94.40     91.81
+[…]
 3 Experiments
-3.1 Cross-Lingual Semantic Discrimination (CLSD)
+3.1 Cross-Lingual Semantic Discrimination
 We propose a task that aims to evaluate multilingual embedding methods based on their capacity to find the closest cross-lingual semantic match amidst large bilingual text collections.
 The goal is to measure the model’s ability to identify a parallel sentence as the most similar to the original sentence in the source language.
-To accomplish this task, a dataset of parallel sentences is required that will be enriched with $N$ distractor sentences in the target language.
-Distractors Generation Prompt Can you provide me with four tricky sentences (numbered) that look structurally and lexically similar but don’t have the same meaning.
-The sentences should be within similar topics and share commonalities with the original.
-Answer in {French/German}!
-{Original Sentence in French/German}
-Figure 1: GPT-4 prompt template used for the generation of distractors
+To accomplish this task, a dataset of parallel sentences is required that will be enriched with four distractor sentences in the target language, all of which are aimed to be semantically similar.
 In this paper, we test the hypothesis that LLMs can generate difficult examples that look similar on the surface, but convey different meanings.
-We prompt GPT-4 (gpt-4-0613) (OpenAI, 2023) to monolingually generate four distractors for each original sentence (see Figure 1).
+We prompt GPT-4 (gpt-4-0613) (OpenAI, 2023) to monolingually generate four distractors for each original sentence (see Prompt 1).
 The distractors are requested to be structurally and lexically similar, whilst being semantically dissimilar.
@@
 CLSD Datasets
-We generate and validate (samples of) four datasets, consisting of both language directions of German/French for the WMT19 (Barrault et al., 2019) and WMT21 (Akhbardeh et al., 2021) test sets.
-[^1] The data is available at link.
+We generate four datasets and validate samples from each, consisting of both language directions of German–French for the WMT19 (Barrault et al., 2019) and WMT21 (Akhbardeh et al., 2021) test sets.
 In the 18,928 generated distractors, we observe that the LLM used different modification strategies such as single polarity inversion, entity replacement or even multiple swaps with word-level Jaccard similarity between originals and distractors at ($\mu=0.47$, $\sigma=0.17$).
-Table 4 illustrates samples of distractors sets.
+Table 5 shows distractor samples.
+Distractor Generation Prompt Can you provide me with four tricky sentences (numbered) that look structurally and lexically similar but don’t have the same meaning.
+The sentences should be within similar topics and share commonalities with the original.
+Answer in {German/French}!
+{Original Sentence in German/French}
+Figure 1: Prompt 1: Distractor generation prompt for GPT-4
 CLSD Evaluation
-To evaluate a multilingual embedding model $E$, we assess its ability to discriminate between original parallel sentence pairs and challenging distractors.
-For a given sentence $S$ in the source language and its corresponding translation $T$ in the target language, along with a set of $N$ distractor sentences $D$ in the target language, the model must produce an embedding of $S$ that is more similar to the embedding of $T$ than to that of any distractor $d_{i}$ in $D$.
-To measure performance of embedding models, we use Precision@1 (P@1), defined as the proportion of samples in which the model ranks the true translation as more similar to the source sentence than the four distractors.
-For example, in Table 2, the column WMT19-DE->FR reports the P@1 of identifying the original French target sentence as more similar than its French distractors.
+To evaluate a multilingual embedding model $E$, we assess its ability to discriminate between a source sentence $S$, its parallel target sentence $T$, and a set of distractors $D$.
+For a given source sentence $S$ and its corresponding translation $T$ in the target language, along with a set of four distractor sentences $D$, the model must produce an embedding of $S$ that is more similar to the embedding of $T$ than to that of any distractor $d_{i}\in D$.
+[…]
 Models Evaluated
@@
 (2024) pre-train a multilingual long-context encoder (8192) and derive a bi-encoder by a process similar to multilingual-e5 that also includes contrastive training against hard negatives.
-paraphrase-multilingual-mpnet-base-v2 (M-MPNet) Reimers and Gurevych (2020) use knowledge distillation through cosine loss training on parallel sentences between an English paraphrase trained teacher model (paraphrase-mpnet-base-v2) and a multilingual student model (XLM-RoBERTa (Conneau et al., 2020)) to teach the student to represent sentences closely across multiple languages.
+paraphrase-multilingual-mpnet-base-v2 (M-MP) Reimers and Gurevych (2020) use knowledge distillation through cosine loss training on parallel sentences between an English paraphrase trained teacher model (paraphrase-mpnet-base-v2) and a multilingual student model (XLM-RoBERTa (Conneau et al., 2020)) to teach the student to represent sentences closely across multiple languages.
+Figure 2: Change in cross-lingual cosine similarity between original and distractor sentence pairs with exactly one token swapped, grouped by the part of speech of the differing token.
 LaBSE Language-agnostic BERT Sentence Embedding (Feng et al., 2022) is trained with translation ranking loss and negative samples.
@@
 The results of the models on the CLSD datasets are shown in Table 2.
-Regarding the direct cross-lingual evaluation, the Bitext Mining specialist LaBSE is the best model in this task with an average score of 94.43 (P@1), demonstrating its consistent ability to cross-lingually rank exact parallel text higher than the distractors.
-Another pattern that emerges is that models (M-E5-b/l, M-GTE) that were also trained on English-centric retrieval datasets, such as the MS Macro (Nguyen et al., 2016), perform worse within our stricter task.
-In the lower half of the table, we evaluate the models using English as a pivot language on both sides, employing the M2M 1.2B translation model (Fan et al., 2021).
-Both LaBSE and M-MPNet show a performance decrease of 2.6/2.8 points compared to direct cross-lingual evaluation.
+Regarding the direct cross-lingual evaluation, the Bitext Mining specialist LaBSE is the best model in this task with an average score of 94.43 (P@1), demonstrating its consistent ability to rank the true translation $T$ higher than any distractor $d_{i}\in D$.
+Another pattern that emerges is that models (M-E5-b/l, M-GTE) that were also trained on English-centric retrieval datasets, such as the MS MARCO (Nguyen et al., 2016), perform worse within our stricter task.
+Pivot Evaluation Setup
+For comparison, we also evaluate models in an English-pivot setting.
+For each CLSD dataset, we translate both the source sentence and its candidate set (true translation plus distractors) into English using the M2M 1.2B model (Fan et al., 2021).
+The evaluation is then performed monolingually in English.
+For example, in the WMT19-DE→FR dataset, the German source sentence and all five French candidates are translated into English, and the model must rank the true English translation above its distractors.
+This yields four additional evaluation setups corresponding to the four CLSD datasets.
+In the pivot evaluation, both LaBSE and M-MP show a performance decrease of 2.6/2.8 points compared to direct cross-lingual evaluation.
 In contrast, models with weaker performance in direct cross-lingual show only minor improvements, except the M-E5-b model, which gains 3.7 points.
@@
 3.3 Fine-grained similarity analysis
-[…]
-$\frac{(1-\cos(E(S),E(T)))-(1-\cos(E(S),E(d))}{\|-\nparallel_{\text{ DE }\leftrightarrow\text{ FR}}}$
-Figure 2 presents model results, showing that the M-MPNet model experiences the largest overall shifts in cosine similarity when single-word modifications are made.
+How do minor surface changes affect the semantic similarity between sentence pairs?
+Can linguistic analysis provide a more precise characterisation of the changes?
+To address these questions, we pick original-distractor pairs $(T,d_{i})$ in which the distractor $d_{i}$ differs from $T$ by only a single word.
+This subset includes 523 German pairs (5.5%) and 324 French pairs (3.4%).
+Using Stanza (v 1.8) (Qi et al., 2020) and further validation by a linguist annotator, we assign part-of-speech tags to the differing token.
+We measure the cross-lingual semantic as the difference between $\cos(E(S),E(T))$ and $\cos(E(S),E(d_{i}))$.
+To improve comparability in our analysis across different models, we normalize [^2] The average cosine similarity difference between parallel and unrelated cross-lingual sentence pairs in the original WMT19/21 datasets is the normalization factor in Figure 2.
+this score by dividing with $\parallel$- $\nparallel_{\text{ DE }\leftrightarrow\text{ FR}}$ for the corresponding model.
+Formally, we measure:
+$\frac{(1-\cos(E(S),E(T)))-(1-\cos(E(S),E(d)))}{\|-\nparallel_{\text{ DE }\leftrightarrow\text{ FR}}}$
+Figure 2 presents model results, showing that the M-MP model experiences the largest overall shifts in cosine similarity when single-word modifications are made.
 The column “ANY” reflects changes across all items, facilitating within-model comparisons to determine which linguistically characterized subgroups exhibit more pronounced changes.
-Notably, LaBSE, the best performing model within CLSD, represents these single-word changes with relatively uniform similarity shifts across linguistic categories, unlike the other models.
+Notably, LaBSE, the best performing model within CLSD, represents single-word changes with relatively uniform similarity shifts across linguistic categories, unlike the other models.
 Additionally, we observe that replacing proper nouns consistently decreases semantic similarity more than replacing other parts of speech across all languages and models.
-Figure 2: Cross-lingual cosine similarity change in original-distractor sentence pairs with exactly one token swapped, grouped by the PoS of the differing token.
 Mono- and Cross-Lingual Change Correlations
@@
 By correlating the results of the cross- and monolingual change for all one-word change cases (i.e.
-“ANY”): The correlation between monolingual and cross-lingual changes is strong for M-MPNet with 0.93 closely followed by M-GTE with 0.85 and LaBSE with 0.82.
+“ANY”): The correlation between monolingual and cross-lingual changes is strong for M-MP with 0.93 closely followed by M-GTE with 0.85 and LaBSE with 0.82.
 The M-E5-b/l shows lower correlation with 0.79 and 0.71 respectively.
@@
 For example, M-GTE shows a low correlation of 0.64 for proper nouns between cross- and monolingual evaluations in French, compared to 0.77 for German.
+Lev Sim. $|D_{\text{bin}}|$    M-E5-b M-E5-l M-GTE M-MP LaBSE
+French
+0.9–1.0 1120    25.9% 20.1% 26.6% 30.4% 32.3%
+0.8–0.9 2138    27.1% 23.9% 31.4% 24.2% 28.2%
+0.7–0.8 2951    31.6% 36.5% 26.2% 28.5% 25.0%
+0.6–0.7 2394    12.6% 16.4% 12.7% 13.0% 12.1%
+0.3–0.6 861    2.8% 3.1% 3.1% 3.9% 2.4%
+German
+0.9–1.0 1523    28.4% 28.7% 30.6% 28.9% 40.4%
+0.8–0.9 2534    32.9% 32.5% 33.2% 35.1% 32.4%
+0.7–0.8 2914    25.4% 23.6% 22.7% 22.2% 19.9%
+0.6–0.7 1890    10.3% 11.8% 10.5% 10.3% 5.9%
+0.3–0.6 690    3.0% 3.4% 3.1% 3.6% 1.5%
+Table 3: Successful distractors for each model, grouped by Levenshtein similarity to the original sentence.
+Values show the fraction of successful distractors (%).
+$|D_{\text{bin}}|$ gives the total number of distractors in each bin.
+3.4 Successful Distractors Analysis
+We conduct an analysis that examines the distribution of successful distractors $d_{i}\in D$ with respect to their Levenshtein similarity to the true translation $T$.
+Specifically, we binned the distractors per model according to their similarity score and report the relative percentage per bin.
+Full results are available in Table 3.
+Across all models, the majority of successful distractors fall into high-similarity bins (especially 0.8–0.99), confirming that minimal changes to the true translation $T$ can suffice to mislead the models.
+However, LaBSE stands out with a relatively lower proportion of successful distractors in lower similarity bins, particularly for German.
+This suggests a greater robustness to more substantial perturbations, potentially explaining its superior performance on our task.
 4 Conclusion
-This paper contributes to cross-lingual semantic search by identifying gaps in current evaluation practices and introducing adversarial examples to test multilingual embeddings.
-Our research, on the German-French case study, introduces four new adversarial cross-lingual evaluation datasets for the news domain.
-Our comparative evaluation of direct cross-lingual and pivot through English evaluation highlights the advantage of each method based on the embedding model used.
-The fine-grained analyses demonstrate how generated distractors offer insights into the cross-lingual capabilities of multilingual models and detail how specific small perturbations in sentences can change these models’ produced semantic representations.
-Future work should extend these approaches to other language pairs and refine adversarial techniques to improve semantic search model evaluations.
+This paper advances cross-lingual semantic search by identifying gaps in current evaluation practices and introducing adversarial examples for testing multilingual embeddings.
+In a German–French case study, we contribute four new adversarial cross-lingual evaluation datasets in the news domain.
+Our comparative experiments highlight that direct cross-lingual and pivot-through-English evaluation each have advantages depending on the embedding model.
+Fine-grained analyses further demonstrate how generated distractors reveal differences in the cross-lingual capabilities of multilingual models and show how small perturbations can affect their semantic representations.
+Future work should extend this approach to additional language pairs and explore more refined adversarial generation methods to improve the evaluation of semantic search models.
 Limitations
-We introduce a pipeline that allows cross-lingual evaluation of semantic search between any two languages that have an available set of parallel texts.
-However, we evaluate only a single language pair (of interest) and only a single model for distractor generation, whereas in reality much cheaper methods might suffice.
-Automatic prediction of PoS tags and named entities in our analysis can result in introducing noise within the fine-grained analysis results.
-The performance of the stanza models and our added manual results validation minimises this risk.
+Our pipeline enables cross-lingual evaluation of semantic search for any language pair with available parallel texts.
+However, we evaluate only one language pair and rely on a single LLM for distractor generation, whereas simpler methods might suffice.
+Automatic prediction of part-of-speech tags and named entities in our fine-grained analysis can introduce noise, although manual validation reduces this risk.
 However, the capabilities of current models in the language of interest should be critically assessed.
-Furthermore, it is important to note that the examination of only the few models is due to space constraints; however, it is worth considering that valuable insights can also be gained from smaller models.
-We hope that larger future work could address these limitations and explore this approach more broadly, creating a large scale cross-lingual semantic search dataset parallel in many, including low-resource languages.
+Moreover, the evaluation covers only a few models due to space constraints; nevertheless, smaller models could also yield valuable insights.
+Future work should address these limitations by broadening the range of languages, models, and adversarial generation techniques, ultimately supporting large-scale evaluation across many languages, including low-resource ones.
 Acknowledgments
-This research is conducted under the project Impresso – Media Monitoring of the Past II Beyond Borders: Connecting Historical Newspapers and Radio.
+This research is conducted under the project Impresso – Media Monitoring of the Past II.
+Beyond Borders: Connecting Historical Newspapers and Radio.
 Impresso is a research project funded by the Swiss National Science Foundation (SNSF 213585) and the Luxembourg National Research Fund (17498891).
+We would also like to thank our colleagues Michelle Wastl and Juri Opitz for their support with annotations and further discussions.
 Appendix A Appendix
@@
 Figure 3: Monolingual cosine similarity change in original-distractor sentence pairs with exactly one token swapped, grouped by the part of speech of the differing token.
-Distractor Intra-Distractor Jaccard Sim.
+Distractor InDistr Jaccard Sim.
 Original: PARIS (Reuters) - Die New Yorker Börse schloss am Freitag mit 0,68 % im Minus und beendete die Woche mit einem allgemeinen Rückgang, da sie von enttäuschenden Ergebnissen, einem Wiederanstieg der Coronavirus-Infektionen und geopolitischen Unsicherheiten beeinträchtigt wurde.
@@
 Adv4: Die Beamten werden im Haushalt für 2021 jedoch nicht belastet werden. 0.818
-Table 4: Original-Distractor sets with their Intra-Distractor Jaccard Similarity. The red font indicates modified text.
+Table 5: Original-Distractor sets with their Intra-Distractor Jaccard Similarity. Red font indicates modified text.
 Original Distractor Tricked Model